OpenStudy (anonymous):

AB and CB form a right angle at the point of intersection, B. If the coordinates of A and B are (8, 0) and (2, 1), respectively, what is the equation of CB?

OpenStudy (anonymous):

x + 6y = 8 6x + y = 11 6x – y = 11 x – 6y = -4